A student at a four-year college wants to know if the mean enrollment at four–year colleges is lower than at two–year colleges in the United States. The student conducted two surveys. Of the 35 two–year colleges surveyed, the mean enrollment was 5,068 with a standard deviation of 4,777. Of the 35 four-year colleges surveyed, the mean enrollment was 5,466 with a standard deviation of 8,191. Find the degrees of freedom you’d use for this test (enter 0 if you would use a distribution that doesn’t use degrees of freedom
Answer #1

solution:

n1 = 35

n2 =35

s1 = 4777

s2 =8191

Population variances are equal.

So we have to use pooled variance.

d.f = n1 + n2 – 2

  =35+35-2 =68

to claim the test degree of freedom is used

it used to find tcritical value from that we can claim rejected or not